OpenStudy (anonymous):

5x+6y=15 Determine the missing coordinate in order pair to satisfy the equation (4, )

OpenStudy (anonymous):

put 4 at the place of x and solve for y

OpenStudy (anonymous):

5x +6y =15 5(4) + 6(-5/6) = 15 5(4)+6y =15 20 + (-5) =15 20 +6y =15 6y =15-20 Sol. (4,-5/6) 6y =-5 y =-5/6